The following information relates to the 2020 debt and equity investment transactions of Flint Ltd., a publicly accountable Canadian corporation. All of the investments were acquired for trading purposes and accounted for using the FV-NI model, with all transaction costs being expensed. No investments were held at December 31, 2019, and the company prepares financial statements only annually, each December 31, following IFRS.
1. On February 1, the company purchased Williams Corp. 12% bonds, at par value for $460,000, plus accrued interest. Interest is payable April 1 and October 1.
2. On April 1, semi-annual interest was received on the Williams bonds.
3. On July 1, 9% bonds of Saint Inc. were purchased. These bonds, with a par value of $168,000, were purchased at par plus accrued interest. Interest dates are June 1 and December 1.
4. On August 12, 2,400 shares of Scotia Corp. were acquired at a cost of $58.00 per share. A 1% commission was paid.
5. On September 1, Williams Corp. bonds with a par value of $92,000 were sold at 103.6 plus accrued interest.
6. On September 28, a dividend of $0.46 per share was received on the Scotia Corp. shares.
7. On October 1, semi-annual interest was received on the remaining Williams Corp. bonds.
8. On December 1, semi-annual interest was received on the Saint Inc. bonds.
9. On December 28, a dividend of $0.48 per share was received on the Scotia Corp. shares.
10. On December 31, the following fair values were determined: Williams Corp. bonds 101.50; Saint Inc. bonds 96; and Scotia Corp. shares $59.50.

Collapse question part

Prepare all 2020 journal entries necessary to properly account for the investment in the Williams Corp. bonds.

Date

Account Titles and Explanation

Debit

Credit

Feb. 1

Apr. 1

Sept. 1

Oct1

DEC 31

RECORD accrue interest

record fair value adjustment
Answer #1
Date Account title Debit credit
Feb 1 FV-NI investment 460000
Interest receivable 18400
cash 478400
[Being investment acquired with 4 month interest accrued (1Oct- 1Feb) 460000*.12*4/12= 18400
April 1 cash (460000*.12*6/12) 27600
Interest receivable 18400
Interest revenue 9200
July 1 FV-NI investment 168000
Interest receivable 1260
cash 169260
[Being investment acquired with 4 month interest accrued (1June -1 July ) 168000*.09*1/12= 1260
August 12 FV-NI investment (2400*58) 139200
Investment income or loss (139200*1%) 1392
cash 140592
Sep 1 cash 99912
Interest revenue 4600
FV-NI investment 92000
Investment income or loss 3312

[Being investment sold along with interest accrued for 5 months (1April -1Sep) 92000*12%*5/12].

Investment gain :92000*(103.6-100)/100= 3312

Sep 28 cash (2400*.46) 1104
Dividend revenue 1104
Oct1 Cash 22080
Interest revenue 22080
[Being interest accrued on remaining bonds :[460000-92000]*.12*6/12
December 1 cash (168000*.09*6/12) 7560
Interest revenue ( 6300
Interest receivable 1260
Dec 28 cash (2400*.48) 1152
Dividend revenue 1152
Dec 31 Interest receivable 12300
Interest revenue 12300
Dec31 FV-NI investment/Fair value adjustment 2400
Investment income or loss 2400

working :

Interest accrual :

Company Par value interest rate period of interest accrued from last payment Interest
Williams corp 460000-92000=368000 12% 3 months   (1Oct-31Dec) 368000*.12*3/12= 11040
Saint Inc 168000 9% 1 month      (1Dec- 31Dec) 168000*.09*1/12=1260
Total interest accrued 12300

Fair value adjustment

cost market value
Williams co. 368000 368000*101.5/100=373520
Saint Inc 168000 168000*96/100= 161280
Scotia corp 139200 2400*59.5= 142800
Total 675200 677600

Investment income due to unrealized holding gain = 677600-675200= 2400